OpenStudy (anonymous):

After being rearranged and simplified, which of the following equations could be solved using the quadratic formula? Check all that apply.

OpenStudy (anonymous):

ax^2 + bx +c = 0 It s quadratic Formula : Delta =b^2- 4ac X1,X2=-b^2.\[\sqrt{\Delta}\]/2a